# Poro 34B and the Blessing of Multilinguality

### v2

@@ Abstract
The pretraining of state-of-the-art large language models now requires trillions of words of text, which is orders of magnitude more than available for the vast majority of languages. While including text in more than one language is an obvious way to acquire more pretraining data, multilinguality is often seen as a curse, and most model training efforts continue to focus near-exclusively on individual large languages. We believe that multilinguality can be a blessing and that it should be possible to substantially improve over the capabilities of monolingual models for small languages through multilingual training. In this study, we introduce Poro 34B, a 34 billion parameter model trained for 1 trillion tokens of Finnish, English, and programming languages, and demonstrate that a multilingual training approach can produce a model that not only substantially advances over the capabilities of existing models for Finnish, but also excels in translation and is competitive in its class in generating English and programming languages. We release the model parameters, scripts, and data under open licenses at https://huggingface.co/LumiOpen/Poro-34B.
@@ 1 Introduction
Neural language models based on the transformer architecture (Vaswani et al., 2017) have led to substantial advances in natural language processing. Encoder-only transformer models such as BERT (Devlin et al., 2019) have advanced the state of the art in a broad range of classification tasks, while decoder-only models such as GPT (Radford et al., 2018) have redefined what can be achieved by generative models, opening new areas of study in e.g. prompting and in-context learning. The success of these models is related in substantial part to their scaling properties: training larger models on more data leads to better results and even entirely new capabilities (Brown et al., 2020). Studies refining our understanding of the optimal balance of model size and training steps have increased the demands on data (Hoffmann et al., 2022b), and many recent models optimize further for inference-time efficiency by training smaller models on more data (Sardana & Frankle, 2023).
@@
These developments have introduced increasing demands for textual data, with many recent models pretrained on a trillion tokens or more (e.g. Touvron et al., 2023; Almazrouei et al., 2023b; MosaicML, 2023; Li et al., 2023; Lozhkov et al., 2024; Groeneveld et al., 2024). While such resources can still be assembled from internet crawls for a few of the languages best represented online, for the vast majority of human languages we have already run out of data for training the largest of language models (Joshi et al., 2020; Villalobos et al., 2022). While it is standard to repeat training data, repetition can lead to reduced sample efficiency and degradation of performance (Hernandez et al., 2022): Muennighoff et al. (2024) estimate that the value of repetition starts to diminish rapidly after four epochs and that repetition ceases to add information around 40 epochs. The availability of data is thus currently a limit for monolingual training for all but a few of the highest-resourced languages.
Multilingual training offers one obvious solution for increasing the amount of training data available, and a large number of multilingual transformer models have been introduced (e.g. Conneau et al., 2020; Lin et al., 2022b; Le Scao et al., 2022; Wei et al., 2023). However, despite the intuitive appeal of augmenting training data with texts in other natural languages, multilinguality is frequently seen as a negative – commonly referred to as the curse of multilinguality (Conneau et al., 2020). While there have been studies of the tradeoffs between monolingual and multilingual training (Fujinuma et al., 2022; Chang et al., 2023) as well as efforts to enhance models specifically for multilinguality (Pfeiffer et al., 2022) and to introduce additional language capabilities to existing models (Gogoulou et al., 2023; Kew et al., 2023; Zhao et al., 2024; Ibrahim et al., 2024), state-of-the-art generative models are still frequently trained near-exclusively on large languages such as English, with only limited efforts specifically focusing on optimizing performance for smaller languages. In this study, we explore how to lift data limitations to create state-of-the-art large generative models from scratch for smaller languages, drawing on the emerging understanding of how to make the most of limited data and turn multilinguality from a curse into a blessing. Some key lessons include 1) limited multilinguality instead of a large number of languages (Conneau et al., 2020; Chang et al., 2023) 2) matching scripts (e.g. Latin) (Fujinuma et al., 2022) and 3) matching language families (Pyysalo et al., 2021), 4) incorporating a cross-lingual signal using translation pairs (Anil et al., 2023; Wei et al., 2023), 5) oversampling target language data up to four epochs (Muennighoff et al., 2024) and 6) augmenting natural language with programming language data (Madaan et al., 2022).
We chose to specifically target the Finnish language, which is an interesting case for study as it is a Uralic language with no large close neighbours in its language family, necessitating more distant transfer than e.g. between English and another Germanic language. While the language is natively spoken by under six million people, its resources are still sufficient to consider a monolingual training approach for larger generative models. In a recent study, Luukkonen et al. (2023) combined several web crawls and curated sources of Finnish to create a dataset of approximately 40B tokens and introduced the monolingual FinGPT models, trained from scratch for 300B tokens (8 epochs). The largest of these models show signs of data limitations, with the 8B parameter model outperforming the 13B in benchmarks. We believe it should be possible to overcome these limitations by applying the lessons listed above. While we cannot match language families, we train for four epochs over the Finnish data and augment it with both English and programming language data as well as an explicit cross-lingual signal from translation pairs. We pursue this approach to create Poro 34B, training a 34B parameter model for a total of 1T tokens – 25 times more than the available Finnish data – and evaluate the model in detail on Finnish, English, and programming language tasks. We find that the model not only achieves the goal of substantially advancing over the performance of existing Finnish models, but is also competitive in its class of open models on English and code as well as remarkably strong in translation tasks.
2 Pretraining data
@@
Figure 1: Pretraining data distribution
For pretraining Poro 34B, we rely on datasets that have been previously preprocessed to remove low-quality texts and boilerplate, filter toxic context, and deduplicate repeated texts. We illustrate the pretraining data distribution in Figure 1 and describe the data briefly in the following. Data sources are detailed in Table 5 in the Reproducibility section.
@@
For Finnish pretraining data, we draw on the resources recently introduced by (Luukkonen et al., 2023) for creating the FinGPT model family. We exclude the ePub and Lehdet resources provided by the National Library of Finland for that work as they could not be shared due to copyright limitations, but use the remaining sources of data, totalling to a 32B token monolingual corpus. The majority of the Finnish data originates from web crawls (approx. 84%) complemented with news sources (approx. 2%), Project Lönnrot, the Finnish equivalent of Project Guthenberg copyright-free book corpus (approx. 0.5%), Wikipedia (approx. 0.5%) and Finnish online discussion forum contents from Reddit and Suomi24 (approx. 13%). Following the rule of thumb proposed by Muennighoff et al. (2024), we upsample the 32B tokens of Finnish so that four epochs over the data are made during training. Consequently approx. 13% of the total tokens seen in pretraining are Finnish.
@@
For English pretraining data, we use SlimPajama (Soboleva et al., 2023), a cleaned and deduplicated subset of the RedPajama corpus [^1] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T (Together Computer, 2023), from which we excluded data from the books category due to their copyright status. We supplemented this dataset with the Project Gutenberg public domain books data from the Dolma corpus [^2] https://huggingface.co/datasets/allenai/dolma (Soldaini et al., 2024). We train for one epoch over the 542B tokens of the English data, which thus represents slightly over half of the 1T total training tokens.
@@
To introduce data representing various programming languages (referred to hereinafter as ”code” for short) into our pretraining, we make use of the Starcoder corpus (Li et al., 2023), a processed subset of The Stack corpus [^3] https://huggingface.co/datasets/bigcode/the-stack (Kocetkov et al., 2023). The original corpus consists of 208B tokens, which we oversample 1.5x so that approximately a third of the tokens seen during pretraining represent code.
@@
We introduce a cross-lingual signal into pretraining by including translation examples from OPUS (Tiedemann, 2009) Specifically, we use the English-Finnish examples from the Tatoeba challenge dataset (Tiedemann, 2020) to generate instruction-formatted translation examples. The Tatoeba challenge training data was reformatted into a minimalistic instruction-following format by recasting each English-Finnish translation pair into a document with the following format:
@@
Where {{en}} and {{fi}} are the English and Finnish texts (resp.) of the translation pair. We additionally reverse the translation order (i.e. Finnish to English instead of English to Finnish) for a total of two documents for each sentence pair. No weighting is applied to the approx. 8B tokens of cross-lingual data, which thus represents slightly under 1% of the pretraining tokens.
@@ 3 Methods
Figure 2: Tokenizer fertility comparison (lower is better)
@@ 3.1 Tokenization
We assess the fertility of the tokenizer on the English and Finnish sentences from the devtest portion of the widely used Flores-101 benchmark for machine translation (Goyal et al., 2022), which allows for a degree of cross-lingual comparability. For code, we use an approx. 1M character sample of lines from the Starcoder held-out test data. [^4] We only sample lines with at least 10 alphabetic characters to avoid very short lines. Figure 2 provides a comparison of the fertility of the tokenizer compared to selected reference tokenizers (see Section 4). We find that on this data the new Poro 34B tokenizer has at least broadly comparable fertility to the lowest-scoring tokenizer on each of Finnish, English, and code, as well as the lowest average fertility of the compared tokenizers.
@@ 3.2 Pretraining
We next briefly present the key model and training parameters (detailed in Table 6 in the Reproducibility section) and the pretraining software and configuration. The hardware used to train the model is described in detail in Appendix A.1.
Architecture Poro 34B is a decoder-only model with a parameter count of 34 billion, sharing its architecture with FinGPT (Luukkonen et al., 2023) and BLOOM (Le Scao et al., 2022). It incorporates an extra layer normalization immediately following the input embedding layer for better training stability and uses ALiBi (Press et al., 2021) as its positional encoding method. The model consists of 54 layers with a hidden dimension of 7168 and a total of 56 attention heads.
Training We train to 1T tokens, intentionally exceeding the Chinchilla compute-opimality estimate (Hoffmann et al., 2022a) of approx. 700B tokens for a model of this size, thus gaining inference-time efficiency for the cost of additional compute investment in pretraining (Sardana & Frankle, 2023). We train with a sequence length of 2048 tokens [^5] We acknowledge that this can be considered limiting by today’s standards, but this limitation can be relieved by methods for extending the context length, for example via linear extrapolation (Press et al., 2021) or interpolation (Al-Khateeb et al., 2023). using a cosine learning rate scheduler with a maximum learning rate of 1.5e-4, decaying to a minimum of 1e-5 over 990B tokens, and a linear warmup of 10B tokens. Our global batch size is 2048 samples totaling to 4194304 tokens per each optimization step.
Software Poro 34B was trained on the LUMI supercomputer GPU-partition, which is powered by AMD MI250X-GPUs. The majority of open source frameworks for large language model pretraining are made to be primarily NVIDIA-compatible, and we required scalable AMD-compatible training software. Thus, we adopted the Megatron-DeepSpeed fork [^6] https://github.com/TurkuNLP/Megatron-DeepSpeed introduced by Luukkonen et al. (2023), which has optimized kernels converted from CUDA to be compatible with AMD ROCm, and has been demonstrated to be a viable solution for large model pretraining on LUMI.
Configuration Considering the hardware available and the selected hyperparameters such as batch size, a configuration of 128 nodes was chosen for the training of the model, resulting in a world size of 1024. The training was done using activation checkpointing, a micro batch size of 1, gradient accumulation of 16, and a 3D parallelism strategy of tensor parallel of degree 2, pipeline parallel of degree 4, resulting in a data parallel degree of 128. This allowed total training cycle throughput of 49618 TFLOPs and 174378 tokens/second.
@@ 3.3 Compute cost
Following Groeneveld et al. (2024), we estimate the carbon footprint of our pretraining by multiplying the theoretical upper bound of the total power used by the GPUs when they are utilized at 100% with the carbon intesity factor of LUMI. Taking into account the systems’s power usage effectiveness (PUE) value of 1.04, [^7] https://www.lumi-supercomputer.eu/sustainable-future/ we approximate the total power consumption to be 448MWh. As LUMI is powered by fully renewable electricity, we assume the carbon intensity factor to be 0.888We acknowledge that this assumption can be contested. As Groeneveld et al. (2024) note: ”LUMI is powered entirely by hydroelectric power and some sources (Ubierna et al., 2022) measure the carbon intensity factor of hydroelectric power to be 0.024.” This brings our emissions to a total of 0 $\mathrm{tCO_{2}eq}$. It is important to note that we only take into account power consumption of the GPUs used, as the consumption of the entire node was not logged during training.
4 Evaluation
We thoroughly analyze the capabilities of the model for Finnish, English and code, first briefly reporting perplexity results and then focusing on community-standard benchmarks for evaluating generative models. We then assess the quality of Finnish text generated by the model and finally evaluate the model’s translation capability from English to Finnish (and vice versa). We primarily compare the performance of the model to a selection of similarly-sized general-purpose open source base language models: Llama 33B [^9] Also known as Llama 30B due to a typo; https://github.com/meta-llama/llama/issues/49 (Touvron et al., 2023), MPT 30B (MosaicML, 2023), and Falcon 40B (Almazrouei et al., 2023a). For Finnish Language evaluations we additionally compare to FinGPT 8B and FinGPT  13B (Luukkonen et al., 2023), and for code we compare also to StarCoder base (Li et al., 2023).
4.1 Data and experimental setup
We benchmark the capabilities of the model in Finnish using the FIN-bench [^10] https://github.com/TurkuNLP/FIN-bench dataset  (Luukkonen et al., 2023), which covers a variety of tasks to assess various aspects of model capabilities in Finnish, combining tasks translated and manually corrected from English BIG-bench (Srivastava et al., 2022) with additional Finnish tasks. We evaluate all FIN-bench results in a 3-shot setting using the standard metrics defined for the benchmark. For English evaluations, we use LM Eval Harness (Gao et al., 2023) to evaluate with the following datasets: ARC Challenge (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022a), and Winogrande (Sakaguchi et al., 2019). We selected these evaluations based on their use as English language benchmarks by Beeching et al. (2023) and use an identical testing configuration here. Programming language proficiency is assessed via the Bigcode Evaluation Harness (Ben Allal et al., 2022) with the HumanEval (Chen et al., 2021), and MBPP (Austin et al., 2021) benchmarks, employing the pass@10 metric for evaluation.
To evaluate the quality of Finnish text generation, we prompt the model with short Finnish phrases and use GPT-4 (Achiam et al., 2023) to judge the quality of the text in terms of coherence, grammatical correctness, and use of specialised vocabulary. Finally, to evaluate translation performance, we use both the Flores-101 devtest (Goyal et al., 2022) data as well as the Tatoeba Challenge test sets (Tiedemann, 2020) in an 8-shot setting, following Zhu et al. (2023).
4.2 Perplexity
@@
Table 1: Character-level perplexity for Poro 34B and selected reference models (lower is better)

### v3

@@ Abstract
The pretraining of state-of-the-art large language models now requires trillions of words of text, which is orders of magnitude more than available for the vast majority of languages. While including text in more than one language is an obvious way to acquire more pretraining data, multilinguality is often seen as a curse, and most model training efforts continue to focus near-exclusively on individual large languages. We believe that multilinguality can be a blessing: when the lack of training data is a constraint for effectively training larger models for a target language, augmenting the dataset with other languages can offer a way to improve over the capabilities of monolingual models for that language. In this study, we introduce Poro 34B, a 34 billion parameter model trained for 1 trillion tokens of Finnish, English, and programming languages, and demonstrate that a multilingual training approach can produce a model that substantially advances over the capabilities of existing models for Finnish and excels in translation, while also achieving competitive performance in its class for English and programming languages. We release the model parameters, scripts, and data under open licenses at https://huggingface.co/LumiOpen/Poro-34B.
@@ 1 Introduction
Neural language models based on the transformer architecture Vaswani et al. (2017) have led to substantial advances in natural language processing. Encoder-only transformer models such as BERT Devlin et al. (2019) have advanced the state of the art in a broad range of classification tasks, while decoder-only models such as GPT Radford et al. (2018) have redefined what can be achieved by generative models, opening new areas of study in prompting and in-context learning. The success of these models is related in substantial part to their scaling properties: training larger models on more data leads to better results and even entirely new capabilities (Brown et al., 2020). Studies refining our understanding of the optimal balance of model size and training steps have increased the demands on data (Hoffmann et al., 2022b), and many recent models optimize further for inference-time efficiency by training smaller models on more data (Sardana and Frankle, 2023).
@@
These developments have introduced increasing demands for textual data, with many recent models pretrained on a trillion tokens or more (e.g. Touvron et al., 2023; Almazrouei et al., 2023; MosaicML, 2023; Li et al., 2023; Lozhkov et al., 2024; Groeneveld et al., 2024). While such resources can still be assembled from internet crawls for a few of the languages best represented online, for the vast majority of human languages we have already run out of data for training the largest of language models (Joshi et al., 2020; Villalobos et al., 2022). While it is standard to repeat training data, repetition can lead to reduced sample efficiency and degradation of performance (Hernandez et al., 2022): Muennighoff et al. (2024) estimate that the value of repetition starts to diminish rapidly after four epochs and that repetition ceases to add information around 40 epochs. The availability of data is thus currently a limit for monolingual training for all but a few of the highest-resourced languages.
@@
Figure 1: Pretraining data distribution.
Multilingual training offers one obvious solution for increasing the amount of training data available, and a large number of multilingual transformer models have been introduced (e.g. Conneau et al., 2020; Lin et al., 2022b; Le Scao et al., 2022; Wei et al., 2023). However, despite the intuitive appeal of augmenting training data with texts in other natural languages, multilinguality is frequently seen as a negative – commonly referred to as the curse of multilinguality (Conneau et al., 2020). While there have been studies of the tradeoffs between monolingual and multilingual training (Fujinuma et al., 2022; Chang et al., 2023) as well as efforts to enhance models specifically for multilinguality (Pfeiffer et al., 2022) and to introduce additional language capabilities to existing models (Gogoulou et al., 2023; Kew et al., 2023; Zhao et al., 2024; Ibrahim et al., 2024), state-of-the-art generative models are still frequently trained near-exclusively on large languages such as English, with only limited efforts specifically focusing on optimizing performance for smaller languages. In this study, we explore how to lift data limitations to create state-of-the-art large generative models from scratch for smaller languages, drawing on the understanding emerging in recent studies on how to make the most of limited data and assure that multilinguality is a blessing rather than a curse. Some key lessons from previous work include 1) limited multilinguality instead of a large number of languages (Conneau et al., 2020; Chang et al., 2023) 2) matching scripts (e.g., Latin) (Fujinuma et al., 2022) and 3) matching language families (Pyysalo et al., 2021), 4) incorporating a cross-lingual signal using translation pairs (Anil et al., 2023; Wei et al., 2023), 5) oversampling target language data up to four epochs (Muennighoff et al., 2024) and 6) augmenting natural language with programming language data (Madaan et al., 2022; Aryabumi et al., 2024).
We chose to specifically target the Finnish language, which is an interesting case for study as it is a Uralic language with no large close neighbours in its language family, necessitating more distant transfer than, for example, between English and another Germanic language. While the language is natively spoken by under six million people, its resources are still sufficient to consider a monolingual training approach for larger generative models. In a recent study, Luukkonen et al. (2023) combined several web crawls and curated sources of Finnish to create a dataset of approximately 40B tokens and introduced the monolingual FinGPT models trained from scratch for 300B tokens. With approximately 8 epochs, the repetition of data is expected to show diminishing returns (Muennighoff et al., 2024), and the largest of these models show signs of data limitations, with the 8B parameter model outperforming the 13B in benchmarks. We believe it should be possible to overcome these limitations by applying the lessons listed above. While we cannot match language families, we train for four epochs over the Finnish data and augment it with both English and programming language data as well as an explicit cross-lingual signal from translation pairs. We pursue this approach to create Poro 34B, training a 34B parameter model for a total of 1T tokens – 25 times more than the available Finnish data – and evaluate the model in detail on Finnish, English, and programming language tasks. We find that the model not only achieves the goal of substantially advancing over the performance of existing Finnish models, but is also competitive in its class of open models on English and code as well as remarkably strong in translation tasks.
2 Pretraining data
For pretraining Poro 34B, we rely on datasets that have been previously preprocessed to remove low-quality texts and boilerplate, filter toxic context, and deduplicate repeated texts. We illustrate the pretraining data distribution in Figure 1 and describe the data briefly in the following. Data sources are detailed in Table 4 in the Appendix.
@@ 2 Pretraining data
For Finnish pretraining data, we draw on the resources recently introduced by Luukkonen et al. for creating the FinGPT model family. We exclude the ePub and Lehdet resources provided by the National Library of Finland for that work as they could not be shared due to copyright limitations, but use the remaining sources of data, totalling to a 32B token monolingual corpus. The majority of the Finnish data originates from web crawls (approx. 84%) complemented with news sources (approx. 2%), Project Lönnrot, the Finnish equivalent of Project Gutenberg copyright-free book corpus (approx. 0.5%), Wikipedia (approx. 0.5%) and Finnish online discussion forum contents from Reddit and Suomi24 (approx. 13%). Following the rule of thumb proposed by Muennighoff et al. (2024), we upsample the 32B tokens of Finnish so that four epochs over the data are made during training. Consequently, approximately 13% of the total tokens seen in pretraining are Finnish.
@@
For English pretraining data, we primarily use SlimPajama (Soboleva et al., 2023), a cleaned and deduplicated subset of the RedPajama corpus [^1] https://huggingface.co/datasets/togethercomputer/RedPajama-Data-1T (Together Computer, 2023), from which we excluded data from the books category due to their copyright status. We supplemented this dataset with the Project Gutenberg public domain books data from the Dolma corpus [^2] https://huggingface.co/datasets/allenai/dolma (Soldaini et al., 2024). We train for one epoch over the 542B tokens of the English data, which thus represents slightly over half of the 1T total training tokens.
@@
To introduce data representing various programming languages (referred to hereinafter as “code" for short) into our pretraining, we make use of the Starcoder corpus (Li et al., 2023), a processed subset of The Stack corpus [^3] https://huggingface.co/datasets/bigcode/the-stack (Kocetkov et al., 2023). The original corpus consists of 208B tokens, which we oversample 1.5x so that approximately a third of the tokens seen during pretraining represent code.
@@
We introduce a cross-lingual signal into pretraining by including translation examples from OPUS (Tiedemann, 2009). Specifically, we use the English-Finnish examples from the Tatoeba dataset (Tiedemann, 2020) to generate instruction-formatted translation examples. The Tatoeba training data was reformatted into a minimalistic instruction-following format by recasting each English-Finnish translation pair into a document with the following format:
@@
Where {{en}} and {{fi}} are the English and Finnish texts (resp.) of the translation pair. We additionally reverse the translation order (i.e., Finnish to English instead of English to Finnish) for a total of two documents for each sentence pair. No weighting is applied to the approximately 8B tokens of cross-lingual data, which thus represents slightly under 1% of the pretraining tokens.
@@ 3 Methods
Figure 2: Tokenizer fertility comparison (lower is better).
@@ 3.1 Tokenization
We assess the fertility of the tokenizer on the English and Finnish sentences from the devtest portion of the widely used Flores-101 benchmark for machine translation Goyal et al. (2022), which allows for a degree of cross-lingual comparability. For code, we use an approximately 1M character sample of lines from the Starcoder held-out test data. [^4] We only sample lines with at least 10 alphabetic characters to avoid very short lines. Figure 2 provides a comparison of the fertility of the tokenizer compared to selected reference tokenizers (see Section 4). We find that on this data the new Poro 34B tokenizer has at least broadly comparable fertility to the lowest-scoring tokenizer on each of Finnish, English, and code, as well as the lowest average fertility of the compared tokenizers.
@@ 3.2 Pretraining
We next briefly present the key model and training parameters (detailed in Table 5 in the Appendix A.1) and the pretraining software and configuration.
Architecture Poro 34B is a decoder-only model with a parameter count of 34 billion, sharing its architecture with FinGPT (Luukkonen et al., 2023) and BLOOM (Le Scao et al., 2022). It incorporates layer normalization immediately following the input embedding layer for better training stability and uses ALiBi (Press et al., 2021) as its positional encoding method. The model consists of 54 layers with a hidden dimension of 7168 and a total of 56 attention heads.
Training We train to 1T tokens, intentionally exceeding the Chinchilla compute-opimality estimate (Hoffmann et al., 2022a) of approximately 700B tokens for a model of this size, thus gaining inference-time efficiency for the cost of additional compute investment in pretraining (Sardana and Frankle, 2023). We train with a sequence length of 2048 tokens [^5] We acknowledge that this can be considered limiting by today’s standards, but this limitation can be relieved by methods for extending the context length, for example via linear extrapolation (Press et al., 2021) or interpolation (Al-Khateeb et al., 2023). using a cosine learning rate scheduler with a maximum learning rate of 1.5e-4, decaying to a minimum of 2e-5 over 990B tokens, and a linear warmup of 10B tokens. Our global batch size is 2048 samples totaling to 4194304 tokens per optimization step.
Software Poro 34B was trained on the LUMI supercomputer GPU partition, which is powered by AMD MI250X GPUs. The majority of open source frameworks for large language model pretraining are made to be primarily NVIDIA-compatible, and we required scalable AMD-compatible training software. Thus, we adopted the Megatron-DeepSpeed fork [^6] https://github.com/TurkuNLP/Megatron-DeepSpeed introduced by Luukkonen et al. (2023), which has optimized kernels converted from CUDA to be compatible with AMD ROCm, and has been demonstrated to be a viable solution for large model pretraining on LUMI. The hardware used to train the model is described in detail in Appendix A.3.
Configuration Considering the hardware available and the selected hyperparameters such as batch size, a configuration of 128 nodes was chosen for the training of the model, resulting in a world size of 1024. The training was done using activation checkpointing, a micro batch size of 1, gradient accumulation of 16, and a 3D parallelism strategy of tensor parallel degree 2, pipeline parallel degree 4, resulting in a data parallel degree of 128. This allowed total training cycle throughput of 49618 TFLOPs and 174378 tokens/second.
@@ 3.3 Compute cost
Following Groeneveld et al. (2024), we estimate the carbon footprint of our pretraining by multiplying the theoretical upper bound of the total power used by the GPUs when they are utilized at 100% with the carbon intensity factor of LUMI. Taking into account the systems’s power usage effectiveness (PUE) value of 1.04, [^7] https://www.lumi-supercomputer.eu/sustainable-future/ we approximate the total power consumption to be 448MWh. As LUMI is powered by fully renewable electricity, we assume the carbon intensity factor to be 0.888We acknowledge that this assumption can be contested. As Groeneveld et al. (2024) note: ”LUMI is powered entirely by hydroelectric power and some sources (Ubierna et al., 2022) measure the carbon intensity factor of hydroelectric power to be 0.024.” This brings our emissions to a total of 0 $\mathrm{tCO_{2}eq}$. It is important to note that we only take into account power consumption of the GPUs used, as the consumption of the entire node was not logged during training.
@@
Table 1: Character-level perplexity for Poro 34B and selected reference models (lower is better).
4 Evaluation
We thoroughly analyze the capabilities of the model for Finnish, English and code, first briefly reporting perplexity results and then focusing on community-standard benchmarks for evaluating generative models. We then assess the quality of Finnish text generated by the model and finally evaluate the model’s translation capability from English to Finnish (and vice versa). For comparison, we include results for the state-of-the-art Finnish language models, FinGPT 8B and FinGPT 13B (Luukkonen et al., 2023), and a selection of similarly-sized general-purpose open source base language models trained on broadly comparable numbers of tokens for English [^9] We chose English models of similar size and training token budget rather than state-of-the-art models to more directly assess the effects of our multilingual training setup on performance in English.: Llama 33B (Touvron et al., 2023), MPT 30B (MosaicML, 2023), and Falcon 40B (Almazrouei et al., 2023). We also provide results for StarCoder base (Li et al., 2023) as a reference for performance on code tasks.
4.1 Data and experimental setup
We benchmark the capabilities of the model in Finnish using the FIN-bench [^10] https://github.com/TurkuNLP/FIN-bench dataset  (Luukkonen et al., 2023), which covers a variety of tasks to assess various aspects of model capabilities in Finnish, combining selected tasks translated and manually corrected from English BIG-bench (Srivastava et al., 2022) with additional Finnish tasks. We evaluate all FIN-bench results in a 3-shot setting using the standard metrics defined for the benchmark. For English evaluations, we use LM Eval Harness (Gao et al., 2023) to evaluate with the following datasets: ARC Challenge (Clark et al., 2018), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2021), TruthfulQA (Lin et al., 2022a), and Winogrande (Sakaguchi et al., 2019). We selected these evaluations based on their use as English language benchmarks by Beeching et al. (2023) and use an identical testing configuration here. Programming language proficiency is assessed via the Bigcode Evaluation Harness (Ben Allal et al., 2022) with the HumanEval (Chen et al., 2021), and MBPP (Austin et al., 2021) benchmarks, employing the pass@10 metric for evaluation.
To evaluate the quality of Finnish text generation, we generate responses to the translated MT-Bench questions with few-shot prompting Zheng et al. (2023). We use a few-shot prompt because this benchmark is designed for chat models and we are evaluating base models. Moreover, we want to unlock the Finnish generation capabilities of the English-focused models by providing in-context examples in Finnish. We use GPT-4 Turbo and human judges to assess the quality of the responses. Finally, to evaluate translation performance, we use both the Flores-101 devtest (Goyal et al., 2022) as well as the Tatoeba test sets (Tiedemann, 2020) in an 8-shot setting, following Zhu et al. (2023).
4.2 Perplexity